40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon May 2 12:51:52 2011, kiwamu, Update, LSC, PRMI locking : plan 
    Reply  Mon May 2 13:30:13 2011, kiwamu, Update, SUS, PRMI locking prep: PRM oplev 
    Reply  Mon May 2 18:16:24 2011, kiwamu, Update, LSC, PRMI locking : status update 
    Reply  Tue May 3 03:53:20 2011, kiwamu, Update, LSC, PRMI locking : done PRMI_locking.png
       Reply  Tue May 3 16:47:55 2011, kiwamu, Update, LSC, PRMI locking : details APtable.png
Message ID: 4617     Entry time: Tue May 3 16:47:55 2011     In reply to: 4604
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: PRMI locking : details 

Here are some details about the PRMI locking done last night.

 

(REFL11 installation)

 REFL11 has been installed on the AP table. The RF signal from the RFPD is sent by a heliax cable which has been called ASDD133.

Before the beam goes into the RFPD a HWP and PBS are installed such that we can adjust the amount of light entering to the photo diode.

One thing I didn't like was that I had to introduce a big amount of the light into the PD to get a reasonably big RF signal.

I was trying to look for an RF signal by looking at a spectrum analyzer, then I realized that the RF signal at 11 MHz was quite tiny when the DC_MON was less than 1.5 V.

After I increased the amount of the light up to 1.9 V in DC_MON, which sounds already too much, I then got able to see the 11 MHz signal on the analyzer.

Note that I decreased the amount of the light down to 0.5 V after I finished locking the PRMI.

We should make sure what is going on with the 11 MHz modulation.

APtable.png

 

(Locking)

First I started locking the MIchelson with AS55. The demodulation phase was already somewhat optimized to the I-signal port.

So I decided not to touch the demodulation phase matrix because it may take some times.

After I eliminated electrical offsets in the digital side, I was easily able to lock the Michelson. The control sign was plus.

Then I started playing with the PRC control too. The demodulation phase in REFL11 looked nearly 45 deg although I didn't carefully measure it.

I made a 45 deg rotational matrix to maximize the I-port signal and tried to lock the PRC. Then immediately I was able to lock PRC as well as MICH.

 GAIN_MICH = 100

 GAIN_PRC = 100

Also GAIN_PRC = -100 gave a carrier resonant lock.

The control filters are the same in MICH and PRC. I used my favorite filters as usual.

 FM1 = 1000 : 10

 FM6 = 0.1 : 1

 FM7 = 1 : 50

Somehow I frequently failed to engage the boost filters (i.e. FM6 and FM7) it looks offsets in the control path kicks either BS or PRM.

Quote from #4604

The PRMI has been successfully locked

Details will be posted in the morning. 

 

ELOG V3.1.3-